### v1

@@ Abstract
This paper addresses the evaluation of MIAs on LLMs with partially inferable training sets, under the ex-post hypothesis, which acknowledges inherent distributional biases between members and non-members datasets. We propose and validate algorithms to create “non-biased” and “non-classifiable” datasets for fairer MIA assessment. Experiments using the Gutenberg dataset on OpenLamma and Pythia show that neutralizing known biases alone is insufficient. Our methods produce non-biased ex-post datasets with AUC-ROC scores comparable to those previously obtained on genuinely random datasets, validating our approach. Globally, MIAs yield results close to random, with only one being effective on both random and our datasets, but its performance decreases when bias is removed.
@@
Membership Inference Attack Assessment Bias.
@@ 1 Introduction
The proliferation of Large Language Models (LLMs) has ignited significant legal and ethical debates, particularly concerning copyright infringement. These models often do not document their training data sources, leading to disputes over unauthorized use of copyrighted material. For instance, lawsuits are piling up against OpenAI accused of training ChatGPT using articles and other books without permission  [^1] See, e.g., businessinsider.com, “The copyright lawsuits against OpenAI are piling up as the tech company seeks data to train its AI”, Jun 30, 2024. Similar accusations have been leveled at Meta or Google for allegedly using protected content  [^2] See, e.g., wired.com, “Congress Senate Tech Companies Pay AI Training Data”, July 2, 2024. These issues underscore the societal, economic, and legal implications of LLM training practices.
@@
Objective. From a technical perspective, determining whether a specific document was a member of the training set of a machine learning (ML) model based on the model’s output, is a Membership Inference Attacks (MIA) problem, highlighted in 2017 by Shokri et al. [26]. However, the effectiveness of MIAs in the context of LLMs is subject to debate.
Limits of existing solutions. To determine whether a particular document has been used to train an AI model, MIA rely on overlearning, which results in stronger predictions when applied to training data. While some research shows that MIAs can achieve high accuracy [19, 18], other studies [7, 6] question their validity due to inherent biases in the datasets of members and non-members used for their assessment. For example, biases like time shifts and n-gram overlaps can lead to over-interpretation of results [7]. Additionally, studies indicate that some biases can be exploited to make a “blind” classifier, without model access, more effective than MIAs [6]. This raises doubts about the robustness and practical relevance of current MIA techniques, and recent surveys like [13] point out the lack of rigor and practical relevance of current proposals.
Research question and contributions. This paper aims to address the problem of assessing MIA effectiveness on LLMs which do not disclose their full training datasets but where part of the training dataset can be inferred. The research question we seek to answer is: How can we construct an unbiased dataset for evaluating MIAs on LLMs with partially inferable training sets?
We propose and evaluate two approaches: (1) Creating datasets that are "non-biased" by design with respect to known biases, and (2) Constructing datasets that cannot be classified, ensuring fairer assessment. Our contributions can be summarized as follows:
@@
Outline. Section 2 reviews related work and positions our research. Section 3 synthesizes our hypothesis and the addressed problem. Section 4 presents our proposed solutions and algorithms, detailing ex-post (i.e., a posteriori) construction of unbiased datasets. Section 5 provides a comparative experimental evaluation of the proposed solutions using the Gutenberg dataset. Finally, Section 6 concludes the paper and suggests directions for future research.
@@ 2 Related Work and Positioning
Methods for adapting Membership Inference Attacks (MIAs), originally developed for machine learning classification algorithms [26], have recently been adapted to the context of Large Language Models (LLMs). Baselines technique use likelihood-based metrics such as Loss [31] and Perplexity [2] to distinguish between members (documents which were used for LLM training) and non-members (not used in training). Perplexity, noted here as $\pi$, is defined for a LLM $\mathcal{M}$ and a document $D$ (e.g., a book) comprising a sequence $d_{1},\ldots,d_{|D|}$ of tokens (words) as the average negative log-likelihood of these token sequences: $\pi(\mathcal{M},D)=-\frac{1}{|D|}\sum_{i=1}^{|D|}\log(\mathds{P}[\mathcal{M}(d_{i}|d_{1},\ldots,d_{i-1})]$, where $|D|$ is the number of tokens in $D$, and $\mathds{P}[d_{2}|d_{1}]$ denotes the conditional probability that token $d_{2}$ follows $d_{1}$. The MIA is performed using a threshold classification, considering $D$ as a member of the training set if the $\pi$ score is lower than $\gamma$: $MIA_{\gamma}(\mathcal{M},D)=\mathds{1}[\pi(\mathcal{M},D)<\gamma]$. Variants include using the perplexity per bit of information after compressing $D$: $MIA_{\gamma}(\mathcal{M},D)=\mathds{1}[\pi(\mathcal{M},D)/zlib(D)<\gamma]$, where $zlib(D)$ is the number of bits after compressing $D$ using Zlib. Several studies show that likelihood-based MIAs applied to LLMs are effective, with high AUC-ROC values. For example, [19] reports an AUC-ROC of $0.856$ for books.
Other metrics, such as Min-k%Prob [25], are based on the premise that if the text has been read by the LLM, it appears with a higher probability. Min-k%Prob selects the ktokens of the document with the minimum probabilities returned by the LLM and computes their average log-likelihood. Results show an AUC-ROC of $0.88$ on ChatGPT for a dataset based on sentences from books, some of which are known to have been read by ChatGPT.
Many other variants of MIAs for LLMs have been proposed, also based on likelihood metrics. Neighboring-based MIAs, where the perplexity score is calibrated using either neighboring models $\mathcal{M}^{\prime}$ or neighboring documents $D^{\prime}$. Classification between members and non-members is then obtained by comparing the likelihood of the target model $\mathcal{M}$ and document $D$ with that of the neighboring model $\mathcal{M}^{\prime}$ and document $D$, or the model $\mathcal{M}$ and neighboring documents $D^{\prime}$. Neighboring models hence assume access to a reference model trained on a disjoint data set drawn from a similar distribution, which is often unrealistic [7]. Neighboring documents [9] are more realistic but present slightly lower performance and additional difficulties in correctly setting noise parameters.
In cases where LLMs do not output likelihood information, complementary metrics can be used with acceptable performance penalty. For example, [14] proposes a MIA method called SaMIA, which measures the similarity between input samples from a document $D$ and the rest of the text in $D$ using ROUGE [16]. SaMIA demonstrates an AUC-ROC of $0.64$ on subsets of The Pile dataset.
@@ 3 Problem statement
A Membership Inference Attack $MIA$ on a large language model $\mathcal{M}$ is a binary classification task aiming to determine whether a textual document $D$ is included in the set of documents $\mathcal{D}_{\text{train}}$ used as the training dataset used to build $\mathcal{M}$. The goal of this attack is to design a function $MIA:\mathcal{D}\rightarrow\{0,1\}$ that can ascertain the truth value of $D\in\mathcal{D}_{\text{train}}$ for any document in the document space $\mathcal{D}$.
@@
The training dataset $\mathcal{D}_{\text{train}}$ of $\mathcal{M}$ is partially inferable, i.e., a subset $\mathcal{D}^{known}_{\text{train}}$ of $\mathcal{D}_{\text{train}}$ can be inferred by the attacker. For example, it is know that OpenAI models like GPT-4 have memorized some precise collections of copyright protected books [3].
@@
A subset $\mathcal{D}^{known}_{nm}$ of non-members (i.e., $\mathcal{D}^{known}_{nm}\subset\mathcal{D}\backslash\mathcal{D}_{\text{train}}$) is (obviously) known. Traditionally, such a subset is constructed by considering documents not available at the time the target LLM was released. This creates inherent biases in the ex-post context, where members and non-members are not drawn from the same ransdom distribution.
We address the problem of producing datasets of members $\mathcal{D}^{Nob}_{train}\subset\mathcal{D}^{known}_{train}$ and non-members $\mathcal{D}^{Nob}_{nm}\subset\mathcal{D}^{known}_{nm}$ which aim to minimize bias (hence the name “Nob” for Non-biased), ensuring a reliable evaluation of MIAs on LLMs while satisfying these three hypotheses.
@@ 4 Neutralizing Bias in Ex-Post Dataset Construction
In this section, we present our approach to identifying and mitigating specific bias in the construction of datsets used for the assesment of MIAs. Our methodology operates in two phases: first, addressing bias due to n-gram overlap, which has been shown to significantly affect the assesment of MIAs [7], and second, addressing additionnal biases that go beyond n-gram overlap.
@@ 4.2 Neutralizing N-gram Bias
5:$hist_{train}\leftarrow histogram(\mathcal{D}^{Nob}_{train},\mathcal{D}^{remain}_{train})$$\triangleright$ Compute n-gram overlap histogram
@@
7:for $i=1$ to $n$ do $\triangleright$ Select document minimizing overlap histograms distance
@@
9:     $D\leftarrow\underset{D\in\mathcal{D}_{nm}^{remain}}{\arg\min}(d_{Kolmogorov-Sminrov}(histogram(\{D\}\cup\mathcal{D}^{Nob}_{nm}),hist_{train})$
@@
  • Histogram computation: It then computes the histogram of n-gram overlap between the selected members and remaining ones (line 3) using the function $histo$ (described below). This histogram represents the target overlap distribution that the non-member set should mirror to be indistinguishable from selected members.
  • Greedy construction: Afterwards, the non-member dataset is constructed document by document, in a greedy fashion: at each step (lines 5 to 9), the document that minimizes the Kolmogorov-Smirnov distance (see below) between the n-gram overlap histograms is selected.
@@
The $histo$ function produces histograms of n-gram overlaps. For a given document $D$, which is considered as a sequence of $k$ tokens (such as letters or words), an n-gram is defined as a continuous sequence of $n$ tokens. The overlap of n-grams from a document $D$ with reference to a set of documents $\mathcal{D}_{ref}$ is computed as the percentage of n-grams in $D$ that appear in any document of $\mathcal{D}_{ref}$. The resulting histogram is a vector where each entry $hist[i]$ corresponds to the number of documents in the first dataset that have an n-gram overlap score, relative to the second dataset, equal to (or close to) value $i$ [^4] More formally, value $hist[i]$ of the histogram is such that $hist[i]=z$ if and only if there are $z$ documents in the first dataset with an n-gram overlap score with value between $i$ and $i+1$.
@@ 4.3 Constructing a Non-Classifiable Dataset
Input $\mathcal{D}^{known}_{train}$, $\mathcal{D}^{known}_{nm}$, integer $n$ expected cardinality of the dataset, $(C_{i})_{i\in[1,N]}$ a vector of agnostic classifiers outputing $\mathds{P}[C_{i}(D)]$ the probability of $D$ being a member
@@
7:while $|\mathcal{D}_{train}^{Nob}|<n$ do $\triangleright$ populating members minimizing confidence
8:     $y\leftarrow\underset{D\in\mathcal{D}_{remain}^{train}}{\arg\min}\left\lVert(C_{i}(D)-0.5)_{i\in[1,N]}\right\rVert_{2}$
9:     $\mathcal{D}_{train}^{Nob}\leftarrow\mathcal{D}_{train}^{Nob}\cup\{y\}$
10:     $\mathcal{D}^{remain}_{train}\leftarrow\mathcal{D}^{remain}_{train}\backslash\{y\}$
11:end while
12:$\mathcal{D}_{nm}^{Nob}\leftarrow\emptyset$
13:$\mathcal{D}^{remain}_{nm}\leftarrow\mathcal{D}^{known}_{nm}\backslash\mathcal{D}_{nm}$
14:while $|\mathcal{D}_{nm}^{Nob}|<n$ do $\triangleright$ populating non-members minimizing confidence
15:     $y\leftarrow\underset{D\in\mathcal{D}_{remain}^{nm}}{\arg\min}\left\lVert(C_{i}(D)-0.5)_{i\in[1,N]}\right\rVert_{2}$
16:     $\mathcal{D}_{nm}^{Nob}\leftarrow\mathcal{D}_{nm}^{Nob}\cup\{y\}$
17:     $\mathcal{D}^{remain}_{nm}\leftarrow\mathcal{D}^{remain}_{nm}\backslash\{y\}$
18:end while
Algorithm overview. In Algorithm 2, we consider a vector of $N$ classifiers $(C_{i})_{i\in[1,N]}$, each of which, once trained, assigns a probability in the range $[0,1]$ to indicate the likelihood of a document being a member. The closer to $1$ (respectively, 00), the more confident the classifier is that the document is a member (resp., non-member). The intuition behind our algorithm is to exploit the confidence to ensures that the constructed datasets are as challenging as possible for the classifiers. It is worth noting that other variants of this algorithm have been implemented, balancing the number of false positives, false negatives, true positives, and true negatives in each member/non-member class. The algorithm operates in the two main steps:
@@
  • Confidence minimization: Using the classifiers $(C_{i})_{i\in[1,N]}$ on the left-out members, we then construct $\mathcal{D}_{train}^{Nob}$ (lines 6 to 10) and $\mathcal{D}_{nm}^{Nob}$ (lines 13 to 17) minimizing the overall confidence of the classifiers. Since the further $C_{i}(D)$ is from $0.5$, the more confident $C_{i}$ is in its assessment of document $D$, we consider the vector $(C_{i}(D)-0.5)_{i\in[1,N]}$ representing the confidence of each classifiers. At each step, we add the element $D$ that minimizes the l2-norm of this confidence vector. For instance, considering the construction of considered members: (1) among the members that have neither been selected as “Non-biased” (Nob) nor used to train the classifiers ($D\in\mathcal{D}^{remain}_{train}$), the one that minimizes the l2-norm of the confidence vector (i.e., $\left\lVert(C_{i}(x)-0.5)_{i\in[1,N]}\right\rVert_{2}$) is selected (line 7); (2) the aforementioned element is inserted in the set (line 8); (3) the set of remaining candidates is updated (line 9).

### v2

@@ Abstract
This paper addresses the evaluation of MIAs on LLMs with partially inferable training sets, under the ex-post hypothesis, which acknowledges inherent distributional biases between members and non-members datasets. We propose and validate algorithms to create “non-biased” and “non-classifiable” datasets for fairer MIA assessment. Experiments using the Gutenberg dataset on OpenLLaMA and Pythia show that neutralizing known biases alone is insufficient. Our methods produce non-biased ex-post datasets on which MIAs achieve AUC-ROC scores comparable to those previously obtained on genuinely random datasets, validating our approach. Globally, MIAs yield results close to random, with only one Meta-Classifier-based MIA being effective on both random and our datasets, but its performance decreases when bias is removed.
@@
Membership Inference Attack LLM Assessment Bias.
@@ 1 Introduction
The proliferation of Large Language Models (LLMs) has ignited significant legal and ethical debates, particularly concerning copyright infringement. These models often do not document their training data sources, leading to disputes over unauthorized use of copyrighted material. For instance, lawsuits are piling up against OpenAI accused of training ChatGPT using articles and other books without permission [^1] See, e.g., “The copyright lawsuits against OpenAI are piling up as the tech company seeks data to train its AI”, Jun 30, 2024. Similar accusations have been leveled at Meta or Google for allegedly using protected content [^2] See, e.g., “Congress Senate Tech Companies Pay AI Training Data”, July 2, 2024. These issues underscore the societal, economic and legal implications of LLM training practices.
@@
Objective. From a technical perspective, determining whether a specific document was a member of the training set of a machine learning (ML) model based on the model’s output, is a Membership Inference Attacks (MIAs) problem, highlighted in 2017 by Shokri et al. [26]. However, the effectiveness of MIAs in the context of LLMs is subject to debate.
Limits of existing solutions. To determine whether a particular document has been used to train an AI model, MIAs rely on overfitting, which results in stronger predictions when applied to training data. While some research shows that MIAs can achieve high accuracy [19, 18], other studies [7, 6] question their validity due to inherent biases in the datasets of members and non-members used for their assessment. For example, biases like time shifts and n-gram overlaps can lead to over-interpretation of results [7]. Additionally, studies indicate that some biases can be exploited to make a “blind” classifier, without model access, more effective than MIAs [6]. This raises doubts about the robustness and practical relevance of current MIA techniques, and recent surveys like [13] point out the lack of rigor and practical relevance of current proposals.
Research question and contributions. This paper aims to address the problem of assessing MIA effectiveness on LLMs which do not disclose their full training datasets but where part of the training dataset can be inferred. The research question we seek to answer is: “How can we construct an unbiased dataset for evaluating MIAs on LLMs with partially inferable training sets?”
We propose and evaluate two approaches: (1) Creating datasets that are “non-biased” by design with respect to known biases, and (2) Constructing datasets that cannot be classified, ensuring fairer assessment. Our contributions can be summarized as follows:
@@
Outline. Section 2 reviews related work and positions our research. Section 3 synthesizes our hypotheses and the addressed problem. Section 4 presents our proposed solutions and algorithms, detailing ex-post (i.e., a posteriori) construction of unbiased datasets. Section 5 provides a comparative experimental evaluation of the proposed solutions. Finally, Section 6 concludes the paper and suggests directions for future research.
@@ 2 Related Work and Positioning
MIA techniques, originally developed for machine learning classification algorithms [26], have recently been adapted to the context of LLMs. The baseline technique use likelihood-based metrics such as loss [31] and perplexity [2] to distinguish between members (documents which were used for LLM training) and non-members (not used in training). Several studies show that likelihood-based MIAs applied to LLMs are effective, with high AUC-ROC values. For example, [19] reports an AUC-ROC of $0.856$ for the Gutenberg dataset [23] (also used for evaluation in our paper). Other metrics, such as Min-k%Prob [25], are based on the premise that if the text has been read by the LLM, it is more likely to appear. More precisely, Min-k%Prob selects the ktokens of the document with the minimum probabilities returned by the LLM and computes their average log-likelihood.
Neighboring-based MIAs calibrate the perplexity score using either neighboring models $\mathcal{M}^{\prime}$ or neighboring documents $D^{\prime}$. Classification between members and non-members is then obtained by comparing the likelihood of the target model $\mathcal{M}$ and document $D$ with that of the neighboring model $\mathcal{M}^{\prime}$ and document $D$, or the model $\mathcal{M}$ and neighboring documents $D^{\prime}$. Neighboring models hence assume access to a reference model trained on a disjoint data set drawn from a similar distribution, which is often unrealistic [7]. Neighboring documents [9] are more realistic but present slightly lower performance and additional difficulties in correctly setting noise parameters.
In cases where LLMs do not output likelihood information, complementary metrics can be used with acceptable performance penalty. For example, [14] proposes a MIA method called SaMIA, which measures the similarity between input samples from a document $D$ and the rest of the text in $D$ using ROUGE [16]. SaMIA demonstrates an AUC-ROC of $0.64$ on subsets of The Pile dataset [10].
@@ 3 Problem Statement
A Membership Inference Attack on a large language model $\mathcal{M}$ is a binary classification task aimed at determining whether a specific textual document $D$ was included in the training dataset $\mathcal{D}_{\text{train}}$ used to build $\mathcal{M}$. The goal of this attack is to design a function $MIA:\mathcal{D}\rightarrow\{0,1\}$ that can ascertain the truth value of $D\in\mathcal{D}_{\text{train}}$ for any document in the document space $\mathcal{D}$.
@@
The training dataset $\mathcal{D}_{\text{train}}$ of $\mathcal{M}$ is partially inferable, i.e., a subset $\mathcal{D}^{known}_{\text{train}}$ of $\mathcal{D}_{\text{train}}$ can be inferred by an attacker. For example, it is know that OpenAI models like GPT-4 have memorized some precise collections of copyright protected books [3].
@@
A subset $\mathcal{D}^{known}_{nm}$ of non-members (i.e., $\mathcal{D}^{known}_{nm}\subset\mathcal{D}\backslash\mathcal{D}_{\text{train}}$) is (obviously) known. Traditionally, such a subset is constructed by considering documents not available at the time the target LLM was released. This creates inherent biases in the ex-post context, where members and non-members are not drawn from the same random distribution.
We address the problem of producing datasets of members $\mathcal{D}^{Nob}_{train}\subset\mathcal{D}^{known}_{train}$ and non-members $\mathcal{D}^{Nob}_{nm}\subset\mathcal{D}^{known}_{nm}$ which aim to minimize bias (hence the name “Nob” for Non-biased). These datasets ensure a reliable evaluation of MIAs on LLMs while satisfying these three hypotheses.
@@ 4 Neutralizing Bias in Ex-Post Dataset Construction
In this section, we present our approach to identifying and mitigating specific bias in the construction of datasets used for the assessment of MIAs. Our methodology operates in two phases: first, addressing bias caused by low n-gram overlap, which has been shown to significantly affect the assessment of MIAs [7]; and second, mitigating additional biases that go beyond n-gram overlap.
@@ 4.2 Neutralizing N-gram Bias
5:$distrib_{train}\leftarrow distribution(\mathcal{D}^{Nob}_{train},\mathcal{D}^{remain}_{train})$triangleright Compute n-gram overlap distribution
@@
7:for $i=1$ to $n$ do $\triangleright$ Select document minimizing overlap distributions distance
@@
9:     $D\leftarrow\underset{D\in\mathcal{D}_{nm}^{remain}}{\arg\min}(d_{Kolmogorov-Sminrov}(distribution(\{D\}\cup\mathcal{D}^{Nob}_{nm}),distrib_{train})$
@@
  • Overlap distribution computation: It then computes the distribution of n-gram overlap between the selected members and remaining ones (line 3) using the function $distribution$ (described below). This distribution represents the target overlap distribution that the non-member set should mirror to be indistinguishable from selected members.
  • Greedy construction: Afterwards, the non-member dataset is constructed document by document, in a greedy fashion: at each step (lines 5 to 9), the document that minimizes the Kolmogorov-Smirnov distance (see below) between the n-gram overlap distributions is selected.
@@
The $distribution$ function produces the distribution of n-gram overlap of a dataset with reference to another. For a given document $D$, which is considered as a sequence of $k$ tokens (such as letters or words), an n-gram is defined as a continuous sequence of $n$ tokens. The overlap of n-grams from a document $D$ with reference to a set of documents $\mathcal{D}_{ref}$ is computed as the percentage of n-grams in $D$ that appear in any document of $\mathcal{D}_{ref}$. The resulting distribution is comprised of the overlap scores of each document in the first dataset.
@@ 4.3 Constructing a Non-Classifiable Dataset
Input $\mathcal{D}^{known}_{train}$, $\mathcal{D}^{known}_{nm}$, integer $n$ the size of the output datasets, $(C_{i})_{i\in[1,N]}$ a vector of agnostic classifiers outputing $\mathds{P}[C_{i}(D)]$ the probability of $D$ being a member
@@
7:$\mathcal{D}_{nm}^{Nob}\leftarrow\emptyset$
8:$\mathcal{D}^{remain}_{nm}\leftarrow\mathcal{D}^{known}_{nm}\backslash\mathcal{D}_{nm}$
9:for $i=1$ to $n$ do $\triangleright$ populating members and non-members minimizing confidence
10:     $D_{m}\leftarrow\underset{D_{m}\in\mathcal{D}_{remain}^{train}}{\arg\min}\left\lVert(C_{i}(D_{m})-0.5)_{i\in[1,N]}\right\rVert_{2}$
11:     $D_{nm}\leftarrow\underset{D_{nm}\in\mathcal{D}_{remain}^{nm}}{\arg\min}\left\lVert(C_{i}(D_{nm})-0.5)_{i\in[1,N]}\right\rVert_{2}$
12:     $\mathcal{D}_{train}^{Nob}\leftarrow\mathcal{D}_{train}^{Nob}\cup\{D_{m}\}$
13:     $\mathcal{D}^{remain}_{train}\leftarrow\mathcal{D}^{remain}_{train}\backslash\{D_{m}\}$
14:     $\mathcal{D}_{nm}^{Nob}\leftarrow\mathcal{D}_{nm}^{Nob}\cup\{D_{nm}\}$
15:     $\mathcal{D}^{remain}_{nm}\leftarrow\mathcal{D}^{remain}_{nm}\backslash\{D_{nm}\}$
16:end for
Algorithm overview. In Algorithm 2, we consider a vector of $N$ classifiers $(C_{i})_{i\in[1,N]}$, each of which, once trained, assigns a probability in the range $[0,1]$ to indicate the likelihood of a document being a member. The closer to $1$ (respectively, 00), the more confident the classifier is that the document is a member (resp., non-member). The intuition behind our algorithm is to exploit the confidence to ensures that the constructed datasets are as challenging as possible for the classifiers. It is worth noting that other variants of this algorithm have been implemented, balancing the number of false positives, false negatives, true positives, and true negatives in each member/non-member class. The algorithm operates in two main steps:
@@
  • Confidence minimization: Using the classifiers $(C_{i})_{i\in[1,N]}$ on the left-out members, we then construct $\mathcal{D}_{train}^{Nob}$ and $\mathcal{D}_{nm}^{Nob}$ (lines 8 to 14) minimizing the overall confidence of the classifiers. Since the further $C_{i}(D)$ is from $0.5$, the more confident $C_{i}$ is in its assessment of document $D$, we consider the vector $(C_{i}(D)-0.5)_{i\in[1,N]}$ representing the confidence of each classifiers. At each step, we add the element $D$ that minimizes the l2-norm of this confidence vector. For instance, considering the construction of considered members: (1) among the members that have neither been selected as “Non-biased” (Nob) nor used to train the classifiers ($D\in\mathcal{D}^{remain}_{train}$), the one that minimizes the l2-norm of the confidence vector (i.e., $\left\lVert(C_{i}(x)-0.5)_{i\in[1,N]}\right\rVert_{2}$) is selected (line 9); (2) the aforementioned element is inserted in the set (line 11); (3) the set of remaining candidates is updated (line 12).